# Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenario

### v1

@@ Abstract
The chiral phase transition in QCD-like theories can be Coleman-Weinberg type supercooling. This is possible to see when Beyond the Standard Model (say, axionlike particle or baryogenesis) is coupled to QCD of the Standard Model (called ordinary QCD). Identifying such a supecooling chiral phase transition thus opens a new window to search for new physics associated with the QCD phase transition epoch of the thermal history, to be probed by gravitational wave and primordial black hole productions. We discuss this possibility by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model in hot and/or dense medium, at the mean-field approximation level. We find that the identification essentially follows the scale violation classification: soft-scale breaking and quantum scale anomaly.
@@ I Introduction
First order phase transitions or supercooling phase transitions currently involve broader interests over theoretical particle physics and cosmology Mazumdar:2018dfl . A particular motivation can be seen in cosmological applications to the thermal history of the Universe: the first-order/supercooling phase transition serves as productions for stochastic gravitational wave (GW) background and primordial black hole (PBH) formation Jung:2021mku ; Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Liu:2021svg ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Lewicki:2024ghw ; Flores:2024lng ; Cai:2024nln ; Florentino:2024kkf ; Goncalves:2024vkj ; Banerjee:2024fam ; Shao:2024dxt ; Banerjee:2024cwv ; Wu:2024lrp ; Hashino:2025fse ; Murai:2025hse ; Franciolini:2025ztf ; Kierkla:2025vwp ; Kanemura:2024pae as nonparticle dark matter candidate Carr:2020gox ; Baldes:2023rqv ; Balaji:2024rvo ; Jinno:2023vnr ; Zhang:2025kbu ; Kanemura:2024pae . It has also been argued that the supercooling electroweak or QCD phase transition can serve as a possible initial condition for slow-roll inflation Zhang:2023acu ; Liu:2024xrh , and subsequent particle production Wang:2022ygk ; Wang:2024tda .
@@
One typical modeling to access those cosmological applications would be a scale-symmetry phase transition via the Coleman-Weinberg (CW) mechanism Coleman:1973jx , assuming an appropriate presence of scalar field (presumably allowing coupled to the standard model plasma). All dimensionful parameters are absent at tree level, but quantum loop corrections induce the scale anomaly and generate the CW potential leading to the spontaneous scale-symmetry breaking. Since the CW potential merely includes a logarithmic field dependence, it exponentially suppresses the tunneling rate at low temperatures. This intrinsic feature naturally induces strong enough supercooling, hence is particularly relevant for the PBH formation Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Baldes:2023rqv ; Conaci:2024tlc ; Arteaga:2024vde ; Salvio:2023ynn ; Salvio:2023blb ; Banerjee:2024cwv ; Kierkla:2025vwp .
@@
Very recently, the importance of what is called the “soft scale-breaking” mass term has also been emphasized in addressing the CW-type supercooling phase transition Zhang:2025kbu . This soft-scale breaking term can dramatically alter the prediction for PBH formation, due to violation of the exponential nucleation approximation. However, it still leaves the successful stochastic GW background predictions essentially as they stand in the case of the pure CW scenario Jaeckel:2016jlh ; Marzola:2017jzl ; Iso:2017uuu ; Ghorbani:2017lyk ; Baldes:2018emh ; Prokopec:2018tnq ; Brdar:2018num ; Marzo:2018nov ; DelleRose:2019pgi ; Ghoshal:2020vud ; Levi:2022bzt ; Kierkla:2022odc ; Salvio:2023qgb ; Zhang:2024vpp . No matter how small such a deformation is, the soft-scale breaking is still active even at low enough temperature – much lower than the critical temperature for the thermal phase transition ($T_{c}$), but as low as the critical temperature for the bubble nucleation ($T_{n}$). Thus this deformation possibility has provided a new testable link between PBH dark matter and GW signatures in the CW-type scenario.
@@
Actually, the potential form Eq.(1) is also the one that arises in low-energy chiral effective models of QCD and QCD-like theories, such as a Nambu-Jona-Lasinio (NJL) description in the mean-field approximation (MFA) with the bosonization of chiral quark bilinear: $M\sim\bar{q}q$. In there, the logarithmic term $\sim M^{4}\ln M$ is generated as a consequence of the quantum scale anomaly, which is also related to the renormalization group evolution of the dynamically induced quartic coupling for $M$. In medium with, e.g., finite baryon chemical potential $\mu_{B}$, this term would get corrections depending on $\mu_{B}$. The soft-scale breaking term $\sim m_{0}^{2}M^{2}$, at $T\sim T_{n}\ll T_{c}$, would be governed by the vacuum contribution, which induces $m_{0}^{2}<0$ as the dynamical chiral symmetry breaking, and also Debye screening mass in medium, $\sim\mu_{B}^{2}M^{2}$, which drives $m_{0}^{2}$ to be positive. Furthermore, beyond the Standard Model contribution, if coupled to the QCD chiral phase transition, might also come in the soft-scale breaking term. Thus, the chiral phase transition in QCD-like theories can also be argued in terms of the same class of the scale violation, the soft-scale breaking $\sim M^{2}$ and the quantum scale anomaly $\sim M^{4}\ln M$, as discussed in Zhang:2025kbu .
In QCD of the Standard Model, which we shall call ordinary QCD, the chiral phase transition goes like crossover at around vanishing baryon chemical potential Aoki:2006we ; Borsanyi:2020fev ; Steinbrecher:2018phh . Thereby, a first-order and supercooling phase transition would be unlikely to take place within ordinary QCD framework. This feature is also reflected in the MFA-NJL description (see, e.g., reviews Hatsuda:1994pi ). The deviation from the CW-type phase transition is tightly linked with the fact that ordinary QCD generates a negatively too large soft-scale breaking mass ($m_{0}^{2}<0$ in Eq.(1)), as will be more explicitized later.
In the literature Shao:2024dxt ; Shao:2024ygm , the possibility of realizing the supercooling from the QCD phase transition has been explored based on a couple of low-energy effective models. There it has been clarified that with a baryon chemical potential much higher than the typical QCD scale $(\lesssim 200-400\,{\rm MeV})$, a small phase transition rate $(\beta/H)$ enough to produce nano hertz GWs can be achieved. This, in terms of the soft-scale breaking, occurs due to not merely the creation of the Fermi surface, but the Debye mass $\sim\mu_{B}^{2}M^{2}$ that is positive and still large enough even at $T\sim T_{n}$, and also.
@@
In this paper, we attempt to clarify how the chiral phase transition in QCD-like theories can be CW-type with Beyond the Standard Model. Our investigation urges us to propose three scenarios: i) many-flavor (walking) dark QCD; ii) QCD with an ALP; iii) the chiral phase transition with a large $\mu_{B}$ $\sim 1$ GeV, where, as done in Shao:2024dxt ; Shao:2024ygm , the Affeck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax is assumed to leave a large amount of $\mu_{B}$ in the QCD phase transition epoch. In all those cases, the scale violation classification works well as well as the CW-type scenario.
@@
This paper is structured as follows: in Sec. II, we present the preliminary setup of the two-flavor NJL model and derive the thermodynamic potential in the MFA with nonzero $\mu_{5},\mu_{B},\theta$ at finite $T$. Section III clarifies the generic properties of the thermal chiral phase transition at $\mu_{B}=0$ in comparison with the CW-type phase transition, particularly paying our attention to the scale violation classes. Then, we propose two classes of “projection” from the chiral phase transition fom in ordinary QCD case into the CW supercooling form: walking dark QCD (Sec. III.1) and QCD with ALP (Sec. III.2). In Sec. IV, we discuss the case with finite $\mu_{B}$, and demonstrate that a large $\mu_{B}$ ($\sim 1$ GeV) can produce the Debye mass properly to make the chiral phase transition CW-typed. The conclusion of this paper is provided in Sec. V.
@@ II Two-flavor mean-field NJL monitor
[…]
We work in the MFA and take into account the mean field only for the $SU(2)$ isospin singlet #1#1#1This assumption is ensured by absence of the isospin chemical potential. Our current discussions can also be straightforwardly extended by including the isospin breaking.. Then the targeted mean fields are those for the following bilinears: $\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)$ and $\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)$. Those are expanded around the mean fields $S^{\prime}=\langle\bar{q}^{\prime}_{i}q^{\prime}_{i}\rangle$ and $P^{\prime}=\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle$ as $\bar{q}^{\prime}_{i}q^{\prime}_{i}=S^{\prime}+\left(:\bar{q}^{\prime}_{i}q^{\prime}_{i}:\right)$ and $\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle=P^{\prime}+\left(:\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}:\right)$2#2#2 The U(1) axial rotation relates the scalar and pseudoscalar bilinears between the original- and prime-base scalar and pseudoscalar bilinears (for each quark flavor $i$) as $\left(\bar{q}_{i}q_{i}\right)$ $=\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}\,,$ $\left(\bar{q}_{i}i\gamma_{5}q_{i}\right)$ $=-\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}\,.$ Here the terms sandwiched by ”:” stand for the normal order product, which means $\langle:\mathcal{O}:\rangle=0$ for $\mathcal{O}=S^{\prime},P^{\prime}$, hence ${\cal L}={\cal L}|_{\rm MF-only}+(:{\cal L}:)$. Thus, in the MFA, the Lagrangian in Eq.(II) takes the form
@@
As is evident in the form of Eq.(6), $\mu_{5}$ and $\theta$ merely act as a catalyzer against the CW-type scale anomaly, i.e., the diving force for the phase transition in the case without them. For instance, even for a large $\mu_{5}$ at $T=\mu_{B}=0$, in the case with massive quarks including physical point, $\mu_{5}$ does not generate the Debye mass or logarithmic term for $M$, hence $\mu_{5}$ itself does not trigger the chiral restoration. In the chiral limit, however, $\mu_{5}$ can act like $\mu_{B}$ as seen from Eq.(7), and therefore can trigger the first order phase transition as has been discussed in the literature Chernodub:2011fr . Even at physical point, a dynamically induced $\mu_{5}$ as the mean field of the axialvector field can also serve as the source for triggering the first order phase transition Shao:2022oqw , which we will not consider in the present study. In what follows, we will thus take $\theta$ and $\mu_{5}$ into account as an option.
III Exploring CW-type chiral phase transition: case with $\mu_{B}=0$
@@
It is well known in the two-flavor NJL model that at the physical point for quark masses, the chiral phase transition at vanishing baryon chemical potentials is crossover (see, e.g., reviews Hatsuda:1994pi ). This is also true even with nonzero $\theta$, as long as $\theta\neq\pi$, at which case the CP phase transition along $P=S^{\prime}$ undergoes like second order Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Sasaki:2013ewa ; Lu:2018ukl ; Huang:2024nbd . Now we understand this phase transition property in terms of the scale violation classes.
@@
Consider first the vacuum case with $T=0$. The tree-level, bare, $(M-m)^{2}$ term has come from the four-fermion interactions along with $g_{s}$ and $g_{d}$, including the shift by the current quark mass. This is a soft-scale breaking term. In addition, the first term in the loop correction part (inside the square bracket in Eq.(9)) includes quadratic $\sim M^{2}\Lambda^{2}$ and logarithmic divergent terms $\sim M^{4}\ln(\Lambda/M)$, where $\Lambda$ denotes the loop momentum cutoff. The former definitely takes negative sign, and gives a destructive contribution to the bare $M^{2}$ term, which destabilizes the chiral symmetric phase at $M=0$ up to the linear shift by $m$, hence triggers the dynamical chiral symmetry breaking. This is another soft-scale breaking term. The net soft-breaking term takes largely negative sign at $T=0$. Meanwhile, the latter logarithmic divergent term takes nothing but a form of the CW-type scale anomaly. Thus, at $T=0$, the effective potential derived from Eq.(9) (with renormalization prescribed) can be cast into essentially the same form as in the generic CW-type potential given in Eq.(1), with a large negative curvature at around $M=0$ ($m_{0}^{2}<0$ in Eq.(1)):
@@
where $a_{f}=2\ln\pi-2\gamma_{E}\simeq 1.14$. As seen from Eqs.(10) and (11), the original CW-type logarithmic term $\sim M^{4}\ln M$ is precisely canceled each other. Thereby, the CW-type first-order phase transition is not realized: no barrier is generated. This is completely consistent with the well-known fact that the two-flavor chiral phase transition in the MFA-NJL model, with $\mu_{B}=0$, is second order in the chiral limit with $m=0$, and crossover at physical point. This phenomenon can also be understood in a view of the original scale anomaly cancellation in an extremely hot medium, which implies a free quark gas picture simply following the Stephan-Boltzmann form of pressure, $\Omega\sim p\sim T^{4}$ (corresponding to the first term in Eq.(11)).
@@
Below we give two possible scenarios with some Beyond the Standard Model to achieve the CW-type chiral phase transition characterized by the near criticality soft-scale breaking in Eq.(12).
III.1 Walking dark QCD

### v2

# Axionlike particle-assisted supercooling chiral phase transition in QCD: Identifying Coleman-Weinberg type-chiral phase transition in QCD-like scenarios

@@ Abstract
We propose a new scenario to realize the Coleman-Weinberg (CW) type chiral phase transition in the QCD thermal history. This scenario predicts an axionlike particle (ALP) with mass $\gtrsim$ sub GeV, consistently with the current experimental and cosmological bounds. The chiral phase transition is evaluated by monitoring ordinary QCD setup in a view of a two-flavor Nambu-Jona-Lasinio model at the mean-field approximation level. The present work thus can open a new window to search for the ALP associated with the QCD phase transition epoch of the thermal history, to be probed by gravitational wave and primordial black hole productions. This proposal is based on a generic classification of the order of the chiral phase transition in a view of the scale violation classes: the soft-scale breaking term and the CW-type scale anomaly term, in or off medium with or without chemical potentials. On this theoretical ground, we also revisit existing scenarios which undergo the supercooling chiral phase transition, such as nearly scale-invariant QCD and QCD with a large baryon chemical potential.
@@ I Introduction
First order phase transitions or supercooling phase transitions currently involve broader interests over theoretical particle physics and cosmology Mazumdar:2018dfl . A particular motivation can be seen in cosmological applications to the thermal history of the Universe: the first-order/supercooling phase transition serves as productions for stochastic gravitational wave (GW) background and primordial black hole (PBH) formation Jung:2021mku ; Kodama:1982sf ; Hall:1989hr ; Kusenko:2020pcg ; Hashino:2021qoq ; Hashino:2022tcs ; He:2022amv ; Liu:2021svg ; Kawana:2022olo ; Lewicki:2023ioy ; Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Salvio:2023blb ; Conaci:2024tlc ; Banerjee:2023brn ; Banerjee:2023vst ; Banerjee:2023qya ; Lewicki:2024ghw ; Flores:2024lng ; Cai:2024nln ; Florentino:2024kkf ; Goncalves:2024vkj ; Banerjee:2024fam ; Shao:2024dxt ; Banerjee:2024cwv ; Wu:2024lrp ; Hashino:2025fse ; Murai:2025hse ; Franciolini:2025ztf ; Kierkla:2025vwp ; Kanemura:2024pae as nonparticle dark matter candidate Carr:2020gox ; Baldes:2023rqv ; Balaji:2024rvo ; Jinno:2023vnr ; Zhang:2025kbu ; Kanemura:2024pae . It has also been argued that the supercooling electroweak or QCD phase transition can provide a possible initial condition for slow-roll inflation Zhang:2023acu ; Liu:2024xrh , and subsequent particle production Wang:2022ygk ; Wang:2024tda .
@@
One typical modeling to access those cosmological applications would be a scale-symmetry phase transition via the Coleman-Weinberg (CW) mechanism Coleman:1973jx , assuming an appropriate presence of scalar field (presumably allowing coupled to the Standard Model (SM) plasma). All dimensionful parameters are absent at tree level, but quantum loop corrections induce the scale anomaly and generate the CW potential leading to the spontaneous scale-symmetry breaking. Since the CW potential merely includes a logarithmic field dependence, it exponentially suppresses the tunneling rate at low temperatures. This intrinsic feature naturally induces strong enough supercooling, hence is particularly relevant for the PBH formation Gouttenoire:2023naa ; Gouttenoire:2023pxh ; Baldes:2023rqv ; Conaci:2024tlc ; Arteaga:2024vde ; Salvio:2023ynn ; Salvio:2023blb ; Banerjee:2024cwv ; Kierkla:2025vwp .
@@
Very recently, the importance of what is called the “soft scale-breaking” mass term has also been emphasized in addressing the CW-type supercooling phase transition Zhang:2025kbu . This soft-scale breaking term can dramatically alter the prediction for PBH formation, due to violation of the exponential nucleation approximation. However, it still leaves the successful stochastic GW background predictions essentially as they stand in the case of the pure CW scenario Jaeckel:2016jlh ; Marzola:2017jzl ; Iso:2017uuu ; Ghorbani:2017lyk ; Baldes:2018emh ; Prokopec:2018tnq ; Brdar:2018num ; Marzo:2018nov ; DelleRose:2019pgi ; Ghoshal:2020vud ; Levi:2022bzt ; Kierkla:2022odc ; Salvio:2023qgb ; Zhang:2024vpp . No matter how small such a deformation is, the soft-scale breaking is still active even at low enough temperature – much lower than the critical temperature for the thermal phase transition ($T_{c}$), but as low as the critical temperature for the bubble nucleation ($T_{n}$). Thus this deformation possibility has provided a new testable link between PBH dark matter and GW signatures in the CW-type scenario Zhang:2025kbu .
@@
Actually, the potential form Eq.(1) is also the one that arises in low-energy chiral effective models of QCD and QCD-like theories, such as a Nambu-Jona-Lasinio (NJL) description in the mean-field approximation (MFA) with the bosonization of chiral quark bilinear: $M\sim\bar{q}q$. In there, the logarithmic term $\sim M^{4}\ln M$ is generated as a consequence of the quantum scale anomaly, which is also related to the renormalization group evolution of the dynamically induced quartic coupling for $M$. In medium with, e.g., finite baryon chemical potential $\mu_{B}$, this term would get corrections depending on $\mu_{B}$. The soft-scale breaking term $\sim m_{0}^{2}M^{2}$, at $T\sim T_{n}\ll T_{c}$, would be governed by the vacuum contribution, which induces $m_{0}^{2}<0$ as the dynamical chiral symmetry breaking, and also Debye screening mass in medium, $\sim\mu_{B}^{2}M^{2}$, which drives $m_{0}^{2}$ to be positive. Furthermore, beyond the SM contribution, if coupled to the QCD chiral phase transition, might also come into play as the soft-scale breaking term. Thus, the chiral phase transition in QCD-like theories can also be argued in terms of the same class of the scale violation, the soft-scale breaking $\sim M^{2}$ and the quantum scale anomaly $\sim M^{4}\ln M$, as discussed in Zhang:2025kbu .
In QCD of the SM, which we shall call ordinary QCD, the chiral phase transition goes like crossover at around vanishing baryon chemical potential Aoki:2006we ; Borsanyi:2020fev ; Steinbrecher:2018phh . Thereby, a first-order and supercooling phase transition would be unlikely to take place within ordinary QCD framework. This feature is also reflected in the MFA-NJL description (see, e.g., reviews Hatsuda:1994pi ). The deviation from the CW-type phase transition is tightly linked with the fact that ordinary QCD generates a negatively too large soft-scale breaking mass ($m_{0}^{2}<0$ in Eq.(1)), as will be more explicitized later.
In the literature Shao:2024dxt ; Shao:2024ygm , the possibility of realizing the supercooling from the QCD phase transition has been explored based on a couple of low-energy effective models. There it has been clarified that with a baryon chemical potential much higher than the typical QCD scale $(\lesssim 200-400\,{\rm MeV})$, a small phase transition rate $(\beta/H)$ enough to produce nano hertz GWs can be achieved. This, in terms of the soft-scale breaking, occurs due to not merely the creation of the Fermi surface, but the Debye mass $\sim\mu_{B}^{2}M^{2}$ that is positive and still large enough even at $T\sim T_{n}$.
@@
In this paper, we first recap and clarify how the chiral phase transition in QCD-like theories can be CW-type with Beyond the SM. The key phenomenon necessary to be triggered is a cancellation of the soft-scale breaking terms at around $T=T_{c}$. We discuss an explicit dynamics to realize this key condition, which includes a well-known feature of many-flavor (walking) dark QCD and the recently addressed chiral phase transition with a large $\mu_{B}$ $\sim 1$ GeV in QCD Shao:2024dxt ; Shao:2024ygm , where for the latter, the Affeck-Dine baryogenesis Affleck:1984fy ; Linde:1985gh ; Dine:2003ax is assumed to leave a large amount of $\mu_{B}$ in the QCD phase transition epoch.
@@
[…]
This paper is structured as follows: in Sec. II, we present the preliminary setup of the two-flavor NJL model and derive the thermodynamic potential in the MFA with nonzero $\mu_{5},\mu_{B},\theta$ at finite $T$. In Sec. III, we clarify the generic properties of the thermal chiral phase transition at $\mu_{B}=0$ in comparison with the CW-type phase transition, particularly paying our attention to the scale violation classes, without $\mu_{B}$ (Sec. III.1) and with $\mu_{B}$ (Sec. III.2). Then, in Sec. IV, we propose the ALP-assisted CW-type chiral supercooling and present the phenomenological constraints on the predicted ALP and discuss its experimental and cosmological detectability. The conclusion of this paper is provided in Sec. V.
@@ II Two-flavor mean-field NJL monitor
$+\frac{g_{s}}{2}\sum^{3}_{a=0}\left[\left(\bar{q}^{\prime}\tau_{a}q^{\prime}\right)^{2}+\left(\bar{q}^{\prime}i\gamma_{5}\tau_{a}q^{\prime}\right)^{2}\right]+g_{d}\left(\text{det}\left[\bar{q}^{\prime}\left(1+\gamma_{5}\right)\bar{q}^{\prime}\right]+\text{h.c.}\right)\,.$ (3)
We work in the MFA and take into account the mean field only for the $SU(2)$ isospin singlet #1#1#1This assumption is ensured by absence of the isospin chemical potential. Our current discussions can also be straightforwardly extended by including the isospin breaking.. Then the targeted mean fields are those for the following bilinears: $\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)$ and $\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)$. Those are expanded around the mean fields $S^{\prime}=\langle\bar{q}^{\prime}_{i}q^{\prime}_{i}\rangle$ and $P^{\prime}=\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle$ as $\bar{q}^{\prime}_{i}q^{\prime}_{i}=S^{\prime}+\left(:\bar{q}^{\prime}_{i}q^{\prime}_{i}:\right)$ and $\langle\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\rangle=P^{\prime}+\left(:\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}:\right)$ #2#2#2The U(1) axial rotation relates the scalar and pseudoscalar bilinears between the original- and prime-base scalar and pseudoscalar bilinears (for each quark flavor $i$) as $\left(\bar{q}_{i}q_{i}\right)$ $=\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}\,$, $\left(\bar{q}_{i}i\gamma_{5}q_{i}\right)$ $=-\left(\bar{q}^{\prime}_{i}q^{\prime}_{i}\right)\text{sin}\frac{\theta}{2}+\left(\bar{q}^{\prime}_{i}i\gamma_{5}q^{\prime}_{i}\right)\text{cos}\frac{\theta}{2}\,.$ . Here the terms sandwiched by ”:” stand for the normal order product, which means $\langle:\mathcal{O}:\rangle=0$ for $\mathcal{O}=S^{\prime},P^{\prime}$, hence ${\cal L}={\cal L}|_{\rm MF-only}+(:{\cal L}:)$. Thus, in the MFA, the Lagrangian in Eq.(II) takes the form
@@
As is evident in the form of Eq.(6), $\mu_{5}$ and $\theta$ merely act as a catalyzer against the CW-type scale anomaly, i.e., the driving force for the phase transition in the case without them. For instance, even for a large $\mu_{5}$ at $T=\mu_{B}=0$, in the case with massive quarks including physical point, $\mu_{5}$ does not generate the Debye mass or logarithmic term for $M$, hence $\mu_{5}$ itself does not trigger the chiral restoration #3#3#3In the chiral limit, $\mu_{5}$ can act like $\mu_{B}$ as seen from Eq.(7), and therefore can trigger the first order phase transition as has been discussed in the literature Chernodub:2011fr . Even at physical point, a dynamically induced $\mu_{5}$ as the mean field of the axialvector field can also serve as the source for triggering the first order phase transition Shao:2022oqw , which we will not consider in the present study. In the present work, we will not consider this possibility just for simplicity.. The MFA-NJL with nonzero $\mu_{5}$, however, involves the high regularization scheme dependence Yu:2015hym . In the literature Farias:2016let a proper subtraction scheme has been shown to support the chiral crossover, not the first order, even in the chiral limit, where $\mu_{5}$ acts as a catalyzer to enhance the chiral/axial breaking via efficiently driving the topological charge fluctuation to the axial anomaly. We will come back to this point later on.
III Exploring CW-type chiral phase transition
In this section, we discuss the features of the chiral phase transition based on the thermodynamic potential in Eq.(6) in light of realization of the CW-type phase transition on the basis of the scale violation classification. The cases we shall study are classified into two: with $\mu_{B}=0$ (Sec. III.1) or $\mu_{B}\neq 0$ (Sec. III.2). Our main phenomenological proposal is in the former case, where an ALP is predicted to deform the conventional chiral crossover/phase transition into the CW-type supercooling, which we call the ALP-assisted CW-type chiral phase transition. This particular scenario will separately be addressed in more detail, in the later section, Sec. IV.
III.1 The case with $\mu_{B}=0$
@@
It is well known in the two-flavor NJL model that at the physical point for quark masses, the chiral phase transition at vanishing baryon chemical potentials is crossover (see, e.g., a review Hatsuda:1994pi ). This is also true even with nonzero $\theta$, as long as $\theta\neq\pi$, at which case the CP phase transition along $P=S^{\prime}$ undergoes like second order Boer:2008ct ; Boomsma:2009eh ; Sakai:2011gs ; Sasaki:2013ewa ; Lu:2018ukl ; Huang:2024nbd . Now we understand this phase transition property in terms of the scale violation classes.
@@
Consider first the vacuum case with $T=0$. The tree-level (or bare) mass term, $(M-m)^{2}$, has come from the four-fermion interactions along with $g_{s}$ and $g_{d}$, including the shift by the current quark mass. This is a soft-scale breaking term. In addition, the first term in the loop correction part (inside the square bracket in Eq.(9)) includes quadratic $\sim M^{2}\Lambda^{2}$ and logarithmic divergent terms $\sim M^{4}\ln(\Lambda/M)$, where $\Lambda$ denotes the loop momentum cutoff. The former definitely takes negative sign, and gives a destructive contribution to the bare $M^{2}$ term, which destabilizes the chiral symmetric phase at $M=0$ up to the linear shift by $m$, hence triggers the dynamical chiral symmetry breaking. This is another soft-scale breaking term. The net soft-breaking term takes largely negative sign at $T=0$. Meanwhile, the latter logarithmic divergent term takes nothing but a form of the CW-type scale anomaly. Thus, at $T=0$, the effective potential derived from Eq.(9) (with renormalization prescribed) can be cast into essentially the same form as in the generic CW-type potential given in Eq.(1), with a large negative curvature at around $M=0$ ($m_{0}^{2}<0$ in Eq.(1)):
@@
where $a_{f}=2\ln\pi-2\gamma_{E}\simeq 1.14$. As seen from Eqs.(10) and (11), the original CW-type logarithmic term $\sim M^{4}\ln M$ is precisely canceled each other. Thereby, the CW-type first-order phase transition is not realized: no barrier is generated. This is completely consistent with the well-known fact that the two-flavor chiral phase transition in the MFA-NJL model, with $\mu_{B}=0$, is second order in the chiral limit with $m=0$, and crossover at physical point. This phenomenon can also be understood in a view of the original scale anomaly cancellation in an extremely hot medium, which implies a free quark gas picture simply following the Stephan-Boltzmann form of pressure, $\Omega\sim p\sim T^{4}$ (corresponding to the first term in Eq.(11)). This cancellation is further promoted to work by nonzero $\mu_{5}$, because $\mu_{5}$ can act as a catalyzer for the chiral/axial breaking when a proper regularization scheme is carried out Farias:2016let .